# A Miyaoka-Yau inequality for hyperplane arrangements in $\mathbb{CP}^{n}$

### v2

@@ 1.1 Main result: klt and CY arrangements
Let $(\mathcal{H},\mathbf{a})$ be a weighted hyperplane arrangement. To avoid writing unnecessary indices, we don’t label the elements of the arrangement. Thus, $\mathcal{H}$ is a finite set of pairwise distinct complex hyperplanes $H\subset\mathbb{CP}^{n}$, and $\mathbf{a}$ is a weight vector in $\mathbb{R}^{\mathcal{H}}$ whose components are positive real numbers $a_{H}>0$ indexed by the elements $H\in\mathcal{H}$.
@@
The set of weights $\mathbf{a}\in\mathbb{R}^{\mathcal{H}}$ such that $(\mathcal{H},\mathbf{a})$ is klt and CY is equal to the relative interior $P^{\circ}$ of the matroid polytope $P$ of $\mathcal{H}$ as a subset of the hyperplane $\sum a_{H}=n+1$. It is well-known that, if the arrangement $\mathcal{H}$ is essential and irreducible, [^3] I.e. $\bigcap_{H\in\mathcal{H}}H=\emptyset$ and $\mathcal{H}$ is not isomorphic to a product -as in Definition 2.5- of two non-empty arrangements. then $\dim P=|\mathcal{H}|-1$ and therefore $P^{\circ}$ is non-empty. See [BGW03, Theorem 1.12.9] and Appendix B for a self contained proof.
@@
Theorem 4.29 asserts that the parabolic bundle $\mathcal{E}_{*}$ on the resolution $X$ is locally abelian. The proof relies on the fact that a collection of divisors $D_{L}$ with $L\in\mathcal{S}\subset\mathcal{L}_{\textup{irr}}$ has non-empty common intersection $\bigcap_{L\in\mathcal{S}}D_{L}$ if and only if the set $\mathcal{S}$ is nested relative to $\mathcal{L}_{\textup{irr}}$ (Definition 3.11). The locally abelian property of $\mathcal{E}_{*}$ is used to define the parabolic second Chern character $\operatorname{par-ch}_{2}(\mathcal{E}_{*})\in H^{4}(X,\mathbb{R})$ (Definition 4.17).
@@ Acknowledgments
We thank Nikolai Mnëv, Artie Prendergast-Smith, and Calum Spicer for useful discussions. We also thank Jorge Vitório Pereira for giving us Example 5.8.
@@ 2.2 Irreducible arrangements and subspaces
A subspace $L\in\mathcal{L}$ of codimension $2$ is irreducible if and only if its multiplicity $m_{L}$ is $\geq 3$.
@@ 2.3 Weighted arrangements
As a direct consequence of Lemma 2.25 , We obtain $3$ slightly different but equivalent formulations of the klt condition.
@@ 3.1.2 Nested subsets of $\mathcal{L}_{\textup{irr}}$ and intersections of divisors $D_{L}$
If $\mathcal{S}$ is nested relative to $\mathcal{L}_{\textup{irr}}$ and $L_{1},\ldots,L_{k}\in\mathcal{S}$ are pairwise non-comparable, then the intersections $\bigcap_{i\in I}L_{i}$ are non-empty and reducible for any subset $I\subset[k]$ with $|I|\geq 2$.
@@ 3.3 Polarization
This is standard and can be proved inductively by considering a single blowup along a smooth subvariety, see item (ii) of Corollary 4.1.4 in [Laz17]. ∎
@@
$\operatorname{vol}_{P_{k}}(D_{L})=\begin{cases}f(k^{n-2})&\textup{ if }\operatorname{codim}L\geq 2,\\ k^{n-1}+f(k^{n-2})&\textup{ if }\operatorname{codim}L=1,\end{cases}$ (3.20)
where $f=\sum_{j=0}^{n-2}C_{j}k^{j}$ is a polynomial in $k$ of degree at most n2n-2.
@@
$c_{1}(P_{k})^{n-1}\cdot\pi^{*}h=k^{n-1}+f(k^{n-2})$
where $f(k^{n-2})$ is as in Lemma 3.40 .
@@ 4.1 General theory
If $k=1$ and $D_{J}=D_{i}$ , then (4.6) is satisfied by taking complements, locally on $D_{i}$, to the subbundle $F^{i}_{<a}\subset F^{i}_{a}$; so that $F^{i}_{a}=F^{i}_{<a}\oplus U_{a}$ with $U_{a}\cong\operatorname{Gr}^{i}_{a}$.
@@
where $n$ is the dimension of $X$. Similarly, the parabolic degree of the induced parabolic structure $\mathcal{V}_{*}$ is defined as
@@ 4.2 Weights and filtration
Recall that the irreducible components of $D=\pi^{-1}(\mathcal{H})$ are of the form $D_{L}$, where $\pi(D_{L})=L$ and $L$ is a non-empty proper irreducible subspace. For $L\in\mathcal{L}_{\textup{irr}}$ let $\mathcal{E}^{L}$ be the subsheaf of $\mathcal{E}$ generated by sections tangent to $\pi^{*}TL$ when restricted to $D_{L}$. [^6] Alternatively, $\mathcal{E}^{L}$ is the pullback of the subsheaf of $T\mathbb{CP}^{n}$ generated by vector fields tangent to $L$. Set
@@ 5.1 Auxiliary results
This example shows that the inequality (5.8) is sharp.
@@ 5.2 Key estimate
Let $\mathcal{V}\subset T\mathbb{CP}^{n}$ be a distribution and let $\operatorname{Tan}(\mathcal{V})$ be the set of all hyperplanes tangent to $\mathcal{V}$.
@@ 6.1 The quadratic form of an arrangement
The next example shows that, for most arrangements, the quadratic form $Q$ is non-degenerate and indefinite.
@@
Let $\mathcal{H}=\{H_{1},\ldots,H_{N}\}$ be an arrangement in $\mathbb{CP}^{n}$ which is generic, in the sense that no $3$ distinct hyperplanes are linearly dependent. Suppose that the number of hyperplanes $N$ is >n+1>n+1. The generic assumption implies that $\sigma_{i}=0$ for all $i$. Equation (6.5) gives us
@@
$Q=(n+1)\sum_{i<j<k}\left(a_{ij}+a_{ik}+a_{jk}\right)^{2}-(n^{2}-1)\cdot\sum_{i<j}a_{ij}^{2}-2\cdot\left(\sum_{i<j}a_{ij}\right)^{2}\,$, (6.7)
@@ 6.3 The matroid polytope and the semistable cone
We present a geometric description of the set of weights for which an arrangement is stable, involving standard constructions of polytopes associated to matroids. For a more in depth discussion and relations to toric geometry, see [Ale15] .
@@
$\frac{Q}{n+1}=\frac{Q_{1}}{n+1}+\frac{Q_{2}}{n_{2}+1}\,.$
@@
6.4 Relation to GIT and stability of pairs
@@ 6.4.3 Stability of pairs
(i) If $s>n+1$ then the rescaling factor $\lambda=(n+1)/s$ in Equation (6.11) is <1<1. Therefore, the weights $a^{\prime}_{H}=\lambda\cdot a_{H}$ of the rescaled arrangement $(\mathcal{H},\mathbf{a}^{\prime})$ satisfy $a^{\prime}_{H}<a_{H}$ for all $H$. This implies that the weak inequalities $\leq$ in (6.25) become strict inequalities <<; showing that $(\mathcal{H},\mathbf{a}^{\prime})$ is klt.
(ii) This is an immediate consequence of item (2) in [Fuj21, Theorem 1.5]. ∎
@@ 7.2 Hirzebruch arrangements
Let $Q:\mathbb{R}^{N}\to\mathbb{R}$ be a real quadratic form. Suppose that $Q\leq 0$ on an open set $U\subset\mathbb{R}^{N}$ and let $\mathbf{x}\in U$. Then the following holds:
@@
    $Q(\mathbf{x})=0$ if and only if $\mathbf{x}$ is in the kernel of $Q$ ;
@@
    if $Q(\mathbf{x})=0$, then $Q\leq 0$ in all of $\mathbb{R}^{N}$.

### v3

@@ Abstract
00footnotetext: 2020 Mathematics Subject Classification. Primary 14N20, 14J60; Secondary 32S22, 52C35.
@@ 1.1 Main result: klt and CY arrangements
Let $\mathcal{H}$ be a finite set of pairwise distinct complex hyperplanes $H\subset\mathbb{CP}^{n}$, and let $\mathbf{a}$ be a vector in $\mathbb{R}^{\mathcal{H}}$ whose components are positive real numbers $a_{H}>0$ indexed by the elements $H\in\mathcal{H}$. We say that $(\mathcal{H},\mathbf{a})$ is a weighted hyperplane arrangement.
@@
The set of weights $\mathbf{a}\in\mathbb{R}^{\mathcal{H}}$ such that $(\mathcal{H},\mathbf{a})$ is klt and CY is equal to the relative interior $P^{\circ}$ of the matroid polytope $P$ of $\mathcal{H}$ as a subset of the hyperplane $\sum a_{H}=n+1$. It is well-known that, if the arrangement $\mathcal{H}$ is essential and irreducible, [^1] I.e. $\bigcap_{H\in\mathcal{H}}H=\emptyset$ and $\mathcal{H}$ is not isomorphic to a product -as in Definition 2.5- of two non-empty arrangements, see Corollary 2.7. then $\dim P=|\mathcal{H}|-1$ and therefore $P^{\circ}$ is non-empty. See [BGW03, Theorem 1.12.9] and Appendix B for a self contained proof.
Calabi-Yau metrics on log pairs. We explain here why Theorem 1.1 can be viewed as a Miyaoka-Yau type inequality in the setting of hyperplane arrangements, by relating it to the existence of singular Calabi-Yau metrics.
We begin by recalling the smooth case. Let $X$ be a compact $n$-dimensional Calabi-Yau manifold and let $\omega$ be a Ricci-flat Kähler metric on $X$. Then the energy of $\omega$, namely the squared $L^{2}$-norm of its curvature $\int_{X}|\mathrm{Riem}|^{2}$ is, up to a positive constant depending on the dimension, given by the topological quantity $\omega]^{n-2}\cdot c_{2}(X)$. In particular, one obtains the inequality $\omega]^{n-2}\cdot c_{2}(X)\geq 0$, with equality if and only if $\omega$ is flat.
We now turn to log pairs $(X,\Delta)$. To match the context of this paper, we assume that $X=\mathbb{CP}^{n}$ and that $\Delta=\sum a_{i}H_{i}$, where $H_{i}\subset\mathbb{CP}^{n}$ are complex hyperplanes and the coefficients $a_{i}$ lie in the interval $(0,1)$. It follows from Kolodziej’s work [Kol98] that, if the weighted arrangement $(\mathcal{H},\mathbf{a})$ is klt and CY, then there exists a (unique up to scale) weak Ricci-flat Kähler metric $\omega_{\operatorname{RF}}$ on $\mathbb{CP}^{n}$ whose volume form has conical-type singularities along $\mathcal{H}$ (see Proposition C.1 in Appendix C for a precise statement).
The metric $\omega_{\operatorname{RF}}$ is smooth on the complement $\mathbb{CP}^{n}\setminus\mathcal{H}$. By analogy with the smooth case, one might expect that the energy $\int_{\mathbb{CP}^{n}\setminus\mathcal{H}}|\mathrm{Riem}|^{2}$ of $\omega_{\operatorname{RF}}$ should be determined by topological data together with the cone angles, or more precisely by a suitable logarithmic second Chern class $c_{2}(X,\Delta)$ associated to the pair. At present, such a formula is far from being established. The main difficulty lies in the limited regularity of $\omega_{\operatorname{RF}}$ near $\mathcal{H}$; in particular, it is not even known in general whether the above energy integral is finite. When $n=2$ some partial progress can be found in [dBS23].
For log smooth pairs, that is, when $\Delta$ has simple normal crossing support, notions of logarithmic second Chern classes $c_{2}(X,\Delta)$ are well established, and Miyaoka-Yau type inequalities involving these classes have been proved; see for instance [Li21, Theorem 1.13]. Moreover, one can easily check that if the arrangement $\mathcal{H}$ is simple normal crossing, then, up to a negative constant depending on the dimension, our quantity $Q(\mathbf{a})$ coincides with the logarithmic second Chern class defined in [Li21, Equation (4)]. However, for such configurations one always has $Q(\mathbf{a})<0$ if $\dim X>1$; see Example 6.6. Thus, in order to capture flat metrics, it is essential to go beyond the normal crossing setting.
In this paper, following [Pan09], we associate a natural parabolic bundle to the pullback of the tangent bundle on a log resolution of the arrangement. This allows us to invoke Mochizuki’s theory [Moc06], where a refined regularity theory for singular connections is available. In this way, we bypass the analytic difficulties arising from the weak Ricci-flat Kähler metric $\omega_{\operatorname{RF}}$.
Finally, we comment on the sign of our inequality $Q(\mathbf{a})\leq 0$, which is opposite to the classical Miyaoka–Yau inequality $\omega]^{n-2}\cdot c_{2}(X)\geq 0$. This discrepancy is explained by the fact that $Q(\mathbf{a})$ should be viewed, roughly speaking, as the second parabolic Chern character $\operatorname{par-ch}_{2}(\mathcal{E}_{*})$ of the parabolic bundle we consider. On the other hand, if $E$ is a vector bundle with $c_{1}(E)=0$, then $\operatorname{ch}_{2}(E)=-c_{2}(E)$, which accounts for the difference in sign.
@@
Theorem 4.29 asserts that the parabolic bundle $\mathcal{E}_{*}$ on the resolution $X$ is locally abelian. The proof of this theorem is non-trivial and relies on the fact that a collection of divisors $D_{L}$ with $L\in\mathcal{S}\subset\mathcal{L}_{\textup{irr}}$ has non-empty common intersection $\bigcap_{L\in\mathcal{S}}D_{L}$ if and only if the set $\mathcal{S}$ is nested relative to $\mathcal{L}_{\textup{irr}}$ (Definition 3.11). The locally abelian property of $\mathcal{E}_{*}$ is used to define the parabolic second Chern character $\operatorname{par-ch}_{2}(\mathcal{E}_{*})\in H^{4}(X,\mathbb{R})$ (Definition 4.17).
@@ Acknowledgments
We thank Nikolai Mnëv, Artie Prendergast-Smith, and Calum Spicer for useful discussions. We also thank Jorge Vitório Pereira for giving us Example 5.8. We would like to thank the referee for a very careful reading of the manuscript and for providing many helpful comments and suggestions that improved the clarity of the paper.
Recent developments
Conjecture 1.3 has recently been resolved in [dBP25]. As an offshoot, a more elementary proof of Theorem 1.1 was obtained in [dBP26]. However, it is worth noting that the methods developed in the present paper (most notably the stability of the parabolic bundle established in Theorem 5.1) remain fundamental; indeed, they play a central role in the proof of Conjecture 1.3 in [dBP25].
@@ 2.2 Irreducible arrangements and subspaces
A subspace $L\in\mathcal{L}$ of codimension $2$ is irreducible if and only if its multiplicity $m_{L}$ is $\geq 3$, c.f. Example 2.3.
@@ 2.3 Weighted arrangements
Recall from Section 1.1 that the weighted arrangement $(\mathcal{H},\mathbf{a})$ is klt if Equation (2.4) holds for any $L\in\mathcal{L}$. As a direct consequence of Lemma 2.25 , we obtain $3$ slightly different but equivalent formulations of the klt condition.
@@ 3.1.2 Nested subsets of $\mathcal{L}_{\textup{irr}}$ and intersections of divisors $D_{L}$
If $\mathcal{S}$ is nested relative to $\mathcal{L}_{\textup{irr}}$ and $L_{1},\ldots,L_{k}\in\mathcal{S}$ are pairwise non-comparable, then the intersections $\bigcap_{i\in I}L_{i}$ are non-empty and reducible for any subset $I\subset[k]$ with $|I|\geq 2$. Indeed, the elements of $\{L_{i}\,|\,i\in I\}\subset\{L_{1},\ldots,L_{k}\}$ are also pairwise non-comparable, therefore by Definition 3.11 their common intersection $\bigcap_{i\in I}L_{i}$ is non-empy and reducible.
@@ 3.3 Polarization
This can be proved inductively by considering a single blowup along a smooth subvariety, see item (ii) of Corollary 4.1.4 in [Laz17]. There is a standard way to make a choice of coefficients $b_{L}$ so that $P_{k}:=k\cdot\pi^{*}(\mathcal{O}(1))-\sum b_{L}D_{L}$ is ample for all $k\gg 1$, for completeness we recall the argument.
Fix an ample divisor $H$ on $X_{0}:=X$ (in our case $H=\mathcal{O}(1)$ and $X_{0}=\mathbb{CP}^{n}$). Let
$X_{r}\xrightarrow{\pi_{r}}X_{r-1}\xrightarrow{\pi_{r-1}}\cdots\xrightarrow{\pi_{1}}X_{0}$
be a sequence of blow-ups, where $\pi_{i}$ is the blow-up of $X_{i-1}$ along a smooth centre $Z_{i}$, and let $E_{i}\subset X_{r}$ denote the total transform of the exceptional divisor of $\pi_{i}$. Writing $\pi:=\pi_{1}\circ\cdots\circ\pi_{r}$, we let
$H_{r}:=\pi^{*}H-\sum_{i=1}^{r}\varepsilon_{i}E_{i}\,$,
for some positive rational numbers
$0<\varepsilon_{r}\ll\cdots\ll\varepsilon_{1}\ll 1.$
The $\mathbb{Q}$-divisor $H_{r}$ is ample on $X_{r}$ . Let $b_{i}$ and $k_{0}$ be positive integers such that $\varepsilon_{i}=b_{i}/k_{0}$ for all $i$, so
$k_{0}\cdot H_{r}=k_{0}\cdot\pi^{*}H-\sum_{i=1}^{r}b_{i}E_{i}$
is ample on $X_{r}$. Thus
$P_{k}:=(k-k_{0})\cdot\pi^{*}H+k_{0}\cdot H_{r}=k\cdot\pi^{*}H-\sum_{i=1}^{r}b_{i}E_{i}$
is is also ample on $X_{r}$ for any $k\geq k_{0}$. ∎
@@
$\operatorname{vol}_{P_{k}}(D_{L})=\begin{cases}f(k)&\textup{ if }\operatorname{codim}L\geq 2,\\ k^{n-1}+f(k)&\textup{ if }\operatorname{codim}L=1,\end{cases}$ (3.20)
where $f(k)=\sum_{j=0}^{n-2}C_{j}k^{j}$ is a polynomial in $k$ of degree at most n2n-2.
@@
$c_{1}(P_{k})^{n-1}\cdot\pi^{*}h=k^{n-1}+f(k)$
where $f(k)$ is as in Lemma 3.40 .
@@ 4.1 General theory
For an interesting source of examples of parabolic bundles on rank $2$ vector bundles over $\mathbb{CP}^{1}$ which relate to the existence of spherical metrics with conical singularities, see [dBP22].
@@
If $k=1$ and $D_{J}=D_{i}$ , then Equation (4.6) is satisfied by taking holomorphic complements to the subbundle $F^{i}_{<a}\subset F^{i}_{a}$ locally along $D_{i}$ in the analytic topology; so that $F^{i}_{a}=F^{i}_{<a}\oplus U_{a}$ with $U_{a}\cong\operatorname{Gr}^{i}_{a}$.
@@
The filtration of $\mathcal{V}$ by the subsheaves $\mathcal{V}^{i}_{a}$ endow the sheaf $\mathcal{V}$ with a parabolic structure as defined in [Moc06, §3.1.1].
@@
where $n$ is the dimension of $X$. In the above equation we identify the top cohomology group $H^{2n}(X,\mathbb{R})$ with $\mathbb{R}$ via integration on $X$, thus we view the cup product as a non-degenerate bilinear pairing $H^{2}(X,\mathbb{R})\times H^{n-2}(X,\mathbb{R})\to\mathbb{R}$. Similarly, the parabolic degree of the induced parabolic structure $\mathcal{V}_{*}$ is defined as
@@ 4.2 Weights and filtration
Recall that the irreducible components of $D=\pi^{-1}(\mathcal{H})$ are of the form $D_{L}$, where $\pi(D_{L})=L$ and $L$ is a non-empty proper irreducible subspace. For $L\in\mathcal{L}_{\textup{irr}}$ let $\mathcal{E}^{L}$ be the subsheaf of $\mathcal{E}$ generated by local sections tangent to $\pi^{*}TL$ when restricted to $D_{L}$. [^4] Alternatively, $\mathcal{E}^{L}$ is the pullback of the subsheaf of $T\mathbb{CP}^{n}$ generated by vector fields tangent to $L$. Set
@@ 5.1 Auxiliary results
This example shows that the inequality (5.8) is sharp. See also Example A.68 in the appendix for a presentation of this distribution in terms of a multivector field.
@@ 5.2 Key estimate
Let $\mathcal{V}\subset T\mathbb{CP}^{n}$ be a distribution and let $\operatorname{Tan}(\mathcal{V})$ be the set of all hyperplanes tangent to $\mathcal{V}$ (see Definition 5.4).
@@
According to the above definition, if $H\notin\operatorname{Tan}(\mathcal{V})$ then there exists $x\in U$ (the regular set of $\mathcal{V}$) such that $\mathcal{V}_{x}\not\subset T_{x}H$. In particular, the subspaces $\mathcal{V}_{x},T_{x}H\subset T_{x}\mathbb{CP}^{n}$ are transversal. By the openness of transversality, it follows that $\mathcal{V}_{y}$ and $T_{y}H$ are transverse for all points $y$ in a Zariski open subset of $H$.
@@ 6.1 The quadratic form of an arrangement
The next example shows that, for generic arrangements, the quadratic form $Q$ is non-degenerate and indefinite.
@@
Let $\mathcal{H}=\{H_{1},\ldots,H_{N}\}$ be an arrangement in $\mathbb{CP}^{n}$ with $n>1$ which is generic, in the sense that no $3$ distinct hyperplanes are linearly dependent. Suppose that the number of hyperplanes $N$ is >n+1>n+1. The generic assumption implies that $\sigma_{i}=0$ for all $i$. Equation (6.5) gives us
@@
$Q=(n+1)\sum_{i<j<k}\left(a_{ij}+a_{ik}+a_{jk}\right)^{2}-2\cdot(n^{2}-1)\cdot\sum_{i<j}a_{ij}^{2}-2\cdot\left(\sum_{i<j}a_{ij}\right)^{2}\,$, (6.7)
@@ 6.3 The matroid polytope and the semistable cone
We present a geometric description of the set of weights for which an arrangement is stable, involving standard constructions of polytopes associated to matroids (see Section 7.4.1 for basics on matroids). For a more in depth discussion and relations to toric geometry, see [Ale15] .
@@
$\frac{Q}{n+1}=\frac{Q_{1}}{n_{1}+1}+\frac{Q_{2}}{n_{2}+1}\,.$
@@
6.4 Relation to GIT and K-stability of pairs
@@ 6.4.3 Stability of pairs
    If $s=n+1$ then $(\mathcal{H},\mathbf{a})$ is stable if and only if $(\mathbb{CP}^{n},\Delta)$ is klt.
  3. (iii)
@@
(i) If $s>n+1$ then the rescaling factor $\lambda=(n+1)/s$ in Equation (6.11) is <1<1. Therefore, the weights $a^{\prime}_{H}=\lambda\cdot a_{H}$ of the rescaled arrangement $(\mathcal{H},\mathbf{a}^{\prime})$ satisfy $a^{\prime}_{H}<a_{H}$ for all $H$. This implies that the weak inequalities $\leq$ in (6.25) become strict inequalities <<; showing that $(\mathcal{H},\mathbf{a}^{\prime})$ is klt. (ii) This is just by Definition 6.11. (iii) This is an immediate consequence of item (2) in [Fuj21, Theorem 1.5]. ∎
@@
Remark 6.42.
A basic result of Odaka [Oda12] asserts that if $(X,L)$ is a polarized complex projective manifold with $c_{1}(X)\leq 0$ (with polarization $L=K_{X}$ if $c_{1}(X)<0$) then $(X,L)$ is $K$-stable. Actually, Odaka allows $X$ to have (semi)-log-canonical singularities if $c_{1}(X)<0$ and log-terminal singularities if $c_{1}(X)=0$. A logarithmic version of Odaka’s result would suggest that if the weighted arrangement $(\mathcal{H},\mathbf{a})$ is either log canonical of general type or klt CY, then $(\mathbb{CP}^{n},\Delta)$ is $K$-stable.
@@ 7.2 Hirzebruch arrangements
Let $Q:\mathbb{R}^{N}\to\mathbb{R}$ be a real quadratic form. Suppose that $Q\leq 0$ on an open set $U\subset\mathbb{R}^{N}$. Then the following holds:
@@
    $Q(\mathbf{x})=0$ for some $\mathbf{x}\in U$ if and only if $\mathbf{x}$ is in the kernel of $Q$ ;
@@
    if $Q(\mathbf{x})=0$ for some $\mathbf{x}\in U$, then $Q\leq 0$ in all of $\mathbb{R}^{N}$.